### v1

# The Andoni-Naor-Neiman inequalities and isometric embeddability into a $\mathrm{CAT}(0)$ space

Abstract.
Andoni, Naor and Neiman (2018) established a family of quadratic metric inequalities that hold true in every $\mathrm{CAT}(0)$ space. As stated in their paper, this family seems to include all previously used quadratic metric inequalities that hold true in every $\mathrm{CAT}(0)$ space. We prove that there exists a metric space that satisfies all inequalities in this family but does not admit an isometric embedding into any $\mathrm{CAT}(0)$ space. More precisely, we prove that the $6$-point metric spaces constructed by Nina Lebedeva, which are known to admit no isometric embedding into any $\mathrm{CAT}(0)$ space, satisfy all inequalities in this family.
This work was supported in part by JSPS KAKENHI Grant Number JP21K03254.
1. Introduction
A condition for points $x_{1},\ldots,x_{n}$ in a metric space $(X,d_{X})$ is called a quadratic metric inequality on $n$ points if it can be written in the form
$0\leq\sum_{i=1}^{n}\sum_{j=1}^{n}a_{ij}d_{X}(x_{i},x_{j})^{2}$ (1)
by some real $n\times n$ matrix $(a_{ij})$. If (1) holds true for all $x_{1},\ldots,x_{n}\in X$, then we say that the metric space $X$ satisfies the quadratic metric inequality (1). We call a quadratic metric inequality a $\mathrm{CAT}(0)$ quadratic metric inequality if every $\mathrm{CAT}(0)$ space satisfies it. Andoni, Naor and Neiman [2] proved the following theorem.
Theorem 1.1 (Andoni, Naor and Neiman [2]).
An $n$-point metric space $X$ admits an isometric embedding into a $\mathrm{CAT}(0)$ space if and only if $X$ satisfies all $\mathrm{CAT}(0)$ quadratic metric inequalities on $n$ points.
For the original statement of Theorem 1.1 in full generality, see [2, Proposition 3]. To find a characterization of those metric spaces that admit an isometric embedding into a $\mathrm{CAT(0)}$ space is a longstanding open problem stated by Gromov in [4, §15] and [3, Section 1.19+] (see also [1, Section 7], [2, Section 1.4] and [6, Section 1]). Theorem 1.1 tells us that we can answer this problem if we can characterize $\mathrm{CAT}(0)$ quadratic inequalities. To see some recent developments in this direction, we recall the following family of $\mathrm{CAT}(0)$ quadratic metric inequalities on $4$ points.
Definition 1.2.
We say that a metric space $(X,d_{X})$ satisfies the $\boxtimes$-inequalities if we have
$0\leq(1-s)(1-t)d_{X}(x,y)^{2}+s(1-t)d_{X}(y,z)^{2}+std_{X}(z,w)^{2}+(1-s)td_{X}(w,x)^{2}\\ -s(1-s)d_{X}(x,z)^{2}-t(1-t)d_{X}(y,w)^{2}.$
for any $s,t\in[0,1]$ and any $x,y,z,w\in X$.
Gromov [4] and Sturm [5] proved independently that every $\mathrm{CAT}(0)$ space satisfies the $\boxtimes$-inequalities. The name $\boxtimes$-inequalities” is based on a notation used by Gromov [4]. Sturm [5] called these inequalities the weighted quadruple inequalities. The present author [6] proved the following theorem.
Theorem 1.3 ([6]).
If a metric space $X$ satisfies the $\boxtimes$-inequalities, then $X$ satisfies all $\mathrm{CAT}(0)$ quadratic metric inequalities on $5$ points.
It follows from Theorem 1.1 and Theorem 1.3 that a metric space $X$ with $|X|\leq 5$ admits an isometric embedding into a $\mathrm{CAT}(0)$ space if and only if $X$ satisfies the $\boxtimes$-inequalities. On the other hand, Nina Lebedeva constructed $6$-point metric spaces that satisfy the $\boxtimes$-inequalities but do not admit an isometric embedding into any $\mathrm{CAT}(0)$ space (see [1, §7.2]). Therefore, there exist $\mathrm{CAT}(0)$ quadratic metric inequalities on $6$ points that do not follow from the $\boxtimes$-inequalities. However to the best of my knowledge, no such $\mathrm{CAT}(0)$ quadratic metric inequality is known explicitly at this moment.
Toward a characterization of $\mathrm{CAT}(0)$ quadratic metric inequalities, Andoni, Naor and Neiman proposed a general procedure to obtain $\mathrm{CAT}(0)$ quadratic metric inequalities, and presented the following large family of $\mathrm{CAT}(0)$ quadratic metric inequalities (see Lemma 27 and Section 5.1 in [2]). Throughout this paper, we denote by $[n]$ the set $\{1,2,\ldots,n\}$ for each positive integer $n$.
Theorem 1.4 (Andoni, Naor and Neiman [2]).
Fix positive integers $m$ and $n$. Let $c_{1},\ldots,c_{m}$ be positive real numbers. Suppose for every $k\in[m]$, $p_{1}^{k},\ldots,p_{n}^{k},q_{1}^{k},\ldots,q_{n}^{k}$ are positive real numbers that satisfy $\sum_{i=1}^{n}p_{i}^{k}=\sum_{j=1}^{n}q_{j}^{k}=1$. Suppose for every $k\in[m]$, $A_{k}=(a_{ij}^{k})$ and $B_{k}=(b_{ij}^{k})$ are $n\times n$ matrices with nonnegative real entries that satisfy
$\sum_{s=1}^{n}a_{is}^{k}+\sum_{s=1}^{n}b_{sj}^{k}=p_{i}^{k}+q_{j}^{k}$
for any $i,j\in[n]$. Then we have
$\sum_{k=1}^{m}\hskip 5.69054pt\sum_{i,j\in[n]\hskip 2.84526pt\textrm{s.t.}\hskip 2.84526pta_{ij}^{k}+b_{ij}^{k}>0}\frac{c_{k}a_{ij}^{k}b_{ij}^{k}}{a_{ij}^{k}+b_{ij}^{k}}d_{X}(x_{i},x_{j})^{2}\leq\sum_{k=1}^{m}\sum_{i=1}^{n}\sum_{j=1}^{n}c_{k}p_{i}^{k}q_{j}^{k}d_{X}(x_{i},x_{j})^{2}$ (2)
for any $\mathrm{CAT}(0)$ space $(X,d_{X})$ and any points $x_{1},\ldots,x_{n}\in X$.
It is easily seen that many known $\mathrm{CAT}(0)$ quadratic metric inequalities, including the $\boxtimes$-inequalities, are of the form (2). Moreover, as stated in [2, Section 5.1], it seems that all of the previously used $\mathrm{CAT}(0)$ quadratic metric inequalities are of the form (2). So it is natural to ask whether the inequalities of the form (2) capture the totality of $\mathrm{CAT}(0)$ quadratic metric inequalities. In this paper, we answer this question negatively by proving the following theorem, which is our main result.
Theorem 1.5.
There exists a $6$-point metric space $(X,d_{X})$ that satisfies the following conditions:
  1. $\mathrm{(i)}$
    For every positive integer $n$, any points $x_{1},\ldots,x_{n}\in X$ satisfy the inequality
    $\frac{1}{2}\sum_{i,j\in[n]\hskip 2.84526pt\textrm{s.t.}\hskip 2.84526pt\pi_{ij}+\pi_{ji}>0}\frac{\pi_{ij}\pi_{ji}}{\pi_{ij}+\pi_{ji}}d_{X}(x_{i},x_{j})^{2}\leq\sum_{i=1}^{n}\sum_{j=1}^{n}p_{i}q_{j}d_{X}(x_{i},x_{j})^{2}$
    for any nonnegative real numbers $p_{1},\ldots,p_{n},q_{1},\ldots,q_{n}$ and any $n\times n$ matrix $\left(\pi_{ij}\right)$ with nonnegative real entries that satisfy
    $\sum_{k=1}^{n}p_{k}=\sum_{k=1}^{n}q_{k}=1,\quad\sum_{k=1}^{n}\pi_{ik}=p_{i}+q_{i}$
    for every $i\in[n]$.
  2. $\mathrm{(ii)}$
    $X$ does not admit an isometric embedding into any $\mathrm{CAT}(0)$ space.
More precisely, we prove that Lebedeva’s $6$-point metric spaces that we have mentioned above satisfy the condition $\mathrm{(i)}$ in Theorem 1.5. As we will show in Section 2, Theorem 1.5 implies the following corollary.
Corollary 1.6.
There exists a $6$-point metric space $(X,d_{X})$ that satisfies the following conditions:
  • $\mathrm{(i)}$
    For every positive integer $n$, any points $x_{1},\ldots,x_{n}\in X$ satisfy the inequality (2) for any $m$, $c_{1},\ldots,c_{m}$, $\left(p_{i}^{1}\right)_{i\in[n]},\ldots,\left(p_{i}^{m}\right)_{i\in[n]}$, $\left(q_{j}^{1}\right)_{j\in[n]},\ldots,\left(q_{j}^{m}\right)_{j\in[n]}$, $A_{1},\ldots,A_{m}$, $B_{1},\ldots,B_{m}$ as in the statement of Theorem 1.4.
  • $\mathrm{(ii)}$
    $X$ does not admit an isometric embedding into any $\mathrm{CAT}(0)$ space.
1.1. Organization of the paper
This paper is organized as follows. In Section 2, we slightly reformulate and generalize the Andoni-Naor-Neiman inequalities (2), and prove that Theorem 1.5 implies Corollary 1.6. In Section 3, we recall the definition of Lebedeva’s $6$-point metric spaces. In Section 4, we prove that Lebedeva’s $6$-point metric spaces satisfy the condition $\mathrm{(i)}$ in Theorem 1.5, which proves Theorem 1.5.
2. A reformulation of the Andoni-Naor-Neiman inequalities
In this section, we slightly reformulate and generalize the Andoni-Naor-Neiman inequalities (2). We first recall the definition of $\mathrm{CAT}(0)$ space.
Definition 2.1.
A complete metric space $(X,d_{X})$ is called a $\mathrm{CAT}(0)$ space if for any $x,y\in X$ and any $t\in[0,1]$, there exists $z\in X$ that satisfies
$d_{X}(w,z)^{2}\leq(1-t)d_{X}(w,x)^{2}+td_{X}(w,y)^{2}-t(1-t)d_{X}(x,y)^{2}$ (3)
for any $w\in X$.
Remark 2.2.
If $d_{X}$ is a semimetric on $X$, or in other words, if $(X,d_{X})$ satisfies the axioms of metric space except for the requirement that $d_{X}$ satisfies the triangle inequalities, then for any $x,y,w\in X$, the triangle inequality
$d_{X}(x,y)\leq d_{X}(w,x)+d_{X}(w,y)$
holds if and only if the quadratic metric inequality
$t(1-t)d_{X}(x,y)^{2}\leq(1-t)d_{X}(w,x)^{2}+td_{X}(w,y)^{2}$ (4)
holds for any $t\in[0,1]$. We remark that this inequality (4) follows from the inequality (3) in Definition 2.1.
Let $(X,d_{X})$ be a $\mathrm{CAT}(0)$ space, and let $x_{1},\ldots,x_{n}\in X$. Suppose $p_{1},\ldots,p_{n}$ are positive real numbers with $\sum_{i=1}^{n}p_{i}=1$. Then it follows from Definition 2.1 that there exists a unique point $z\in X$ that satisfies
$d_{X}(w,z)^{2}+\sum_{i=1}^{n}p_{i}d_{X}(z,x_{i})^{2}\leq\sum_{i=1}^{n}p_{i}d_{X}(w,x_{i})^{2}$ (5)
for any $w\in X$ (see Lemma 4.4 and Theorem 4.9 in [5]). The point $z\in X$ with the above property is called the barycenter of the probability measure $\mu=\sum_{i=1}^{n}p_{i}\delta_{x_{i}}$ on $X$, and denoted by $\mathrm{bar}(\mu)$, where $\delta_{x_{i}}$ is the Dirac measure at $x_{i}$. It is easily seen that the point $z$ in Definition 2.1 is no other than $\mathrm{bar}\left((1-t)\delta_{x}+t\delta_{y}\right)$.
Although the following proposition and its corollary are just a slight reformulation of Lemma 27 in [2], they will play a key role in the proof of Theorem 1.5. We denote by $\binom{[n]}{2}$ the set of all $2$-element subsets of $[n]$.
Proposition 2.3 (cf. Lemma 27 in [2]).
Fix a positive integer $n$. Suppose $p_{1},\ldots,p_{n}$, $q_{1},\ldots,q_{n}$ are nonnegative real numbers that satisfy $\sum_{i=1}^{n}p_{i}=\sum_{j=1}^{n}q_{j}=1$. Suppose $\left(\pi_{ij}\right)$ is an $n\times n$ matrix with nonnegative real entries that satisfies
$\sum_{j=1}^{n}\pi_{ij}=p_{i}+q_{i}$
for every $i\in[n]$. Suppose $(X,d_{X})$ is a $\mathrm{CAT}(0)$ space and $x_{1},\ldots,x_{n}\in X$. We set
$z=\mathrm{bar}\left(\sum_{i=1}^{n}p_{i}\delta_{x_{i}}\right),\quad z_{ij}=\mathrm{bar}\left(\frac{\pi_{ij}}{\pi_{ij}+\pi_{ji}}\delta_{x_{i}}+\frac{\pi_{ji}}{\pi_{ij}+\pi_{ji}}\delta_{x_{j}}\right)$
for any $i,j\in[n]$ with $\pi_{ij}+\pi_{ji}>0$. Then we have
$\sum_{i,j\in[n]\hskip 2.84526pt\textrm{s.t.}\pi_{ij}+\pi_{ji}>0}\pi_{ij}d_{X}(z,z_{ij})^{2}+\frac{1}{2}\sum_{i,j\in[n]\hskip 2.84526pt\textrm{s.t.}\hskip 2.84526pt\pi_{ij}+\pi_{ji}>0}\frac{\pi_{ij}\pi_{ji}}{\pi_{ij}+\pi_{ji}}d_{X}(x_{i},x_{j})^{2}\leq\sum_{i=1}^{n}\sum_{j=1}^{n}p_{i}q_{j}d_{X}(x_{i},x_{j})^{2}.$
Proof.
It follows from the inequality (5) that
$d_{X}(x_{j},z)^{2}+\sum_{i=1}^{n}p_{i}d_{X}(x_{i},z)^{2}\leq\sum_{i=1}^{n}p_{i}d_{X}(x_{i},x_{j})^{2}$
holds for each $j\in[n]$. By multiplying this inequality by $q_{j}$ and summing over $j\in[n]$, we obtain
$\sum_{j=1}^{n}q_{j}d_{X}(x_{j},z)^{2}+\sum_{i=1}^{n}p_{i}d_{X}(x_{i},z)^{2}\leq\sum_{i=1}^{n}\sum_{j=1}^{n}p_{i}q_{j}d_{X}(x_{i},x_{j})^{2}.$
Since we have $z_{ij}=z_{ji}$ for any $i,j\in[n]$ with $\pi_{ij}+\pi_{ji}>0$, it follows from (3) that
$\sum_{j=1}^{n}$ $q_{j}d_{X}(x_{j},z)^{2}+\sum_{i=1}^{n}p_{i}d_{X}(x_{i},z)^{2}=\sum_{k=1}^{n}(p_{k}+q_{k})d_{X}(x_{k},z)^{2}=\sum_{k=1}^{n}\sum_{l=1}^{n}\pi_{kl}d_{X}(x_{k},z)^{2}$
$=\sum_{k=1}^{n}\pi_{kk}d_{X}(x_{k},z)^{2}+\sum_{\{k,l\}\in\binom{[n]}{2}\hskip 2.84526pt\textrm{s.t.}\hskip 2.84526pt\pi_{kl}+\pi_{lk}>0}(\pi_{kl}+\pi_{lk})\left(\frac{\pi_{kl}}{\pi_{kl}+\pi_{lk}}d_{X}(x_{k},z)^{2}+\frac{\pi_{lk}}{\pi_{kl}+\pi_{lk}}d_{X}(x_{l},z)^{2}\right)$
$\geq\sum_{k\in[n]\hskip 2.84526pt\textrm{s.t.}\pi_{kk}>0}\pi_{kk}d_{X}(z_{kk},z)^{2}+\sum_{\{k,l\}\in\binom{[n]}{2}\hskip 2.84526pt\textrm{s.t.}\hskip 2.84526pt\pi_{kl}+\pi_{lk}>0}(\pi_{kl}+\pi_{lk})\left(d_{X}(z,z_{kl})^{2}+\frac{\pi_{kl}\pi_{lk}}{(\pi_{kl}+\pi_{lk})^{2}}d_{X}(x_{k},x_{l})^{2}\right)$
$=\sum_{i,j\in[n]\hskip 2.84526pt\textrm{s.t.}\pi_{ij}+\pi_{ji}>0}\pi_{ij}d_{X}(z,z_{ij})^{2}+\frac{1}{2}\sum_{i,j\in[n]\hskip 2.84526pt\textrm{s.t.}\hskip 2.84526pt\pi_{ij}+\pi_{ji}>0}\frac{\pi_{ij}\pi_{ji}}{\pi_{ij}+\pi_{ji}}d_{X}(x_{i},x_{j})^{2},$
which proves the proposition. ∎
Remark 2.4.
Because the statement of Proposition 2.3 is symmetric with respect to $(p_{1},\ldots,p_{n})$ and $(q_{1},\ldots,q_{n})$ except for the definition of $z$, the same statement holds true even if we set $z=\mathrm{bar}(\sum_{j=1}^{n}q_{j}\delta_{x_{j}})$ instead of $z=\mathrm{bar}(\sum_{i=1}^{n}p_{i}\delta_{x_{i}})$.
The following corollary follows immediately from Proposition 2.3.
Corollary 2.5 (cf. Lemma 27 in [2]).
Fix a positive integer $n$. Suppose $p_{1},\ldots,p_{n}$, $q_{1},\ldots,q_{n}$, and $(\pi_{ij})$ are as in the statement of Proposition 2.3. Suppose $(X,d_{X})$ is a $\mathrm{CAT}(0)$ space and $x_{1},\ldots,x_{n}\in X$. Then we have
$\frac{1}{2}\sum_{i,j\in[n]\hskip 2.84526pt\textrm{s.t.}\hskip 2.84526pt\pi_{ij}+\pi_{ji}>0}\frac{\pi_{ij}\pi_{ji}}{\pi_{ij}+\pi_{ji}}d_{X}(x_{i},x_{j})^{2}\leq\sum_{i=1}^{n}\sum_{j=1}^{n}p_{i}q_{j}d_{X}(x_{i},x_{j})^{2}.$ (6)
We define the following condition.
Definition 2.6.
Fix a positive integer $n$. Let $(X,d_{X})$ be a semimetric space. We say that $X$ satisfies the $\mathrm{ANN}(n)$ inequalities if any points $x_{1},\ldots,x_{n}\in X$ satisfy the inequality (6) for any $p_{1},\ldots,p_{n},q_{1},\ldots,q_{n}$, and $(\pi_{ij})$ as in the statement of Proposition 2.3.
Remark 2.7.
It is easily seen that for every $t\in[0,1]$, the inequality (4) belongs to the $\mathrm{ANN}(n)$ inequalities for any $n\geq 3$. Therefore, a semimetric space is a metric space whenever it satisfies the $\mathrm{ANN}(n)$ inequalities for some integer $n\geq 3$.
To set up some basic facts about the $\mathrm{ANN}(n)$ inequalities, we recall the following elementary facts, which follows from straightforward computation.
Proposition 2.8.
Suppose $a$, $a_{1}$, $a_{2}$, $b$, $b_{1}$, $b_{2}$ are nonnegative real numbers such that $a=a_{1}+a_{2}>0$ and $b=b_{1}+b_{2}>0$. Then we have
$\frac{a_{1}b_{1}}{a_{1}+b_{1}}+\frac{a_{2}b_{2}}{a_{2}+b_{2}}\leq\frac{ab}{a+b}.$
We will use the following proposition in Section 4.
Proposition 2.9.
Let $n$ be a positive integer, and let $(X=\{x_{1},\ldots,x_{n}\},d_{X})$ be a semimetric space with $|X|=n$. Assume that the inequality
$\frac{1}{2}\sum_{\{i,j\}\in[n]\hskip 2.84526pt\textrm{s.t.}\hskip 2.84526pt\pi_{ij}+\pi_{ji}>0}\frac{\pi_{ij}\pi_{ji}}{\pi_{ij}+\pi_{ji}}d_{X}(x_{i},x_{j})^{2}\leq\sum_{i=1}^{n}\sum_{j=1}^{n}p_{i}q_{j}d_{X}(x_{i},x_{j})^{2}$
holds true for any $p_{1},\ldots,p_{n},q_{1},\ldots,q_{n}$, and $(\pi_{ij})$ as in the statement of Proposition 2.3. Then $X$ satisfies the $\mathrm{ANN}(m)$ inequalities for all positive integers $m$.
Proof.
Fix a positive integer $m$. Fix nonnegative real numbers $p_{1},\ldots,p_{m}$, $q_{1},\ldots,q_{m}$ and an $m\times m$ matrix $(\pi_{ij})$ with nonnegative real entries such that
$\sum_{k=1}^{m}p_{k}=\sum_{k=1}^{m}q_{k}=1,\quad\sum_{k=1}^{m}\pi_{ik}=p_{i}+q_{i}$
for every $i\in[m]$. Choose a map $\varphi:[m]\to[n]$ arbitrarily. For each $k,l\in[n]$, we set
$\tilde{p}_{k}=\sum_{i\in\varphi^{-1}(\{k\})}p_{i},\qquad\tilde{q}_{k}=\sum_{i\in\varphi^{-1}(\{k\})}q_{i},\qquad\tilde{\pi}_{kl}=\sum_{(i,j)\in\varphi^{-1}(\{k\})\times\varphi^{-1}(\{l\})}\pi_{ij}.$
Then it is easily seen that we have
$\sum_{k=1}^{n}\tilde{p}_{k}=\sum_{k=1}^{n}\tilde{q}_{k}=1,\quad\sum_{k=1}^{n}\tilde{\pi}_{ik}=\tilde{p}_{i}+\tilde{q}_{i}$
for every $i\in[n]$. It follows from Proposition 2.8 and the hypothesis that
$\frac{1}{2}$ $\sum_{i,j\in[m]\textrm{ s.t. }\pi_{ij}+\pi_{ji}>0}\frac{\pi_{ij}\pi_{ji}}{\pi_{ij}+\pi_{ji}}d_{X}(x_{\varphi(i)},x_{\varphi(j)})^{2}$
$=\frac{1}{2}\sum_{k,l\in[n]}\left(\sum_{(i,j)\in\varphi^{-1}(\{k\})\times\varphi^{-1}(\{l\})\textrm{ s.t. }\pi_{ij}+\pi_{ji}>0}\frac{\pi_{ij}\pi_{ji}}{\pi_{ij}+\pi_{ji}}\right)d_{X}(x_{k},x_{l})^{2}$
$\leq\frac{1}{2}\sum_{k,l\in[n]\textrm{ s.t. }\tilde{\pi}_{kl}+\tilde{\pi}_{lk}>0}\left(\frac{\tilde{\pi}_{kl}\tilde{\pi}_{lk}}{\tilde{\pi}_{kl}+\tilde{\pi}_{lk}}\right)d_{X}(x_{k},x_{l})^{2}$
$\leq\sum_{k=1}^{n}\sum_{l=1}^{n}\tilde{p}_{k}\tilde{q}_{l}d_{X}(x_{k},x_{l})^{2}$
$=\sum_{k=1}^{n}\sum_{l=1}^{n}\left(\sum_{i\in\varphi^{-1}(\{k\})}p_{i}\right)\left(\sum_{j\in\varphi^{-1}(\{l\})}q_{j}\right)d_{X}(x_{k},x_{l})^{2}$
$=\sum_{k=1}^{n}\sum_{l=1}^{n}\left(\sum_{i\in\varphi^{-1}(\{k\})}\sum_{j\in\varphi^{-1}(\{l\})}p_{i}q_{j}d_{X}(x_{\varphi(i)},x_{\varphi(j)})^{2}\right)$
$=\sum_{i=1}^{m}\sum_{j=1}^{m}p_{i}q_{j}d_{X}(x_{\varphi(i)},x_{\varphi(j)})^{2},$
which proves the proposition. ∎
The following corollary follows from Proposition 2.9 immediately.
Corollary 2.10.
Suppose $m$ and $n$ are positive integers with $m\leq n$, and $X$ is a semimetric space that satisfies the $\mathrm{ANN}(n)$ inequalities. Then $X$ satisfies the $\mathrm{ANN}(m)$ inequalities.
The following lemma clarifies the relation between the $\mathrm{ANN}(n)$ inequalities and the inequalities (2) established by Andoni-Naor-Neiman [2].
Lemma 2.11.
Fix a positive integer $n$ and nonnegative real numbers $p_{1},\ldots,p_{n}$, $q_{1},\ldots,q_{n}$ that satisfy $\sum_{i=1}^{n}p_{i}=\sum_{j=1}^{n}q_{j}=1$. Let $(X,d_{X})$ be a semimetric space. Suppose $x_{1},\ldots,x_{n}\in X$ are points that satisfy the inequality (6) for any $n$ by $n$ matrix $\left(\pi_{ij}\right)$ with nonnegative real entries that satisfies $\sum_{j=1}^{n}\pi_{ij}=p_{i}+q_{i}$ for every $i\in[n]$. Then the inequality
$\sum_{i,j\in[n]\hskip 2.84526pt\textrm{s.t.}\hskip 2.84526pta_{ij}+b_{ij}>0}\frac{a_{ij}b_{ij}}{a_{ij}+b_{ij}}d_{X}(x_{i},x_{j})^{2}\leq\sum_{i=1}^{n}\sum_{j=1}^{n}p_{i}q_{j}d_{X}(x_{i},x_{j})^{2}.$
holds true for any $n$ by $n$ matrices $A=(a_{ij})$ and $B=(b_{ij})$ with nonnegative real entries that satisfy $\sum_{k=1}^{n}a_{ik}+\sum_{k=1}^{n}b_{kj}=p_{i}+q_{j}$ for any $i,j\in[n]$.
Proof.
Fix $n\times n$ matrices $A=(a_{ij})$ and $B=(b_{ij})$ with nonnegative real entries that satisfy
$\sum_{k=1}^{n}a_{ik}+\sum_{k=1}^{n}b_{kj}=p_{i}+q_{j}$
for any $i,j\in[n]$. Set $\pi_{ij}=a_{ij}+b_{ji}$ for any $i,j\in[n]$. Then we have $\sum_{j=1}^{n}\pi_{ij}=p_{i}+q_{i}$ for every $i\in[n]$ clearly. It follows from Proposition 2.8 that
$\frac{a_{ij}b_{ij}}{a_{ij}+b_{ij}}+\frac{a_{ji}b_{ji}}{a_{ji}+b_{ji}}\leq\frac{(a_{ij}+b_{ji})(a_{ji}+b_{ij})}{(a_{ij}+b_{ji})+(a_{ji}+b_{ij})}=\frac{\pi_{ij}\pi_{ji}}{\pi_{ij}+\pi_{ji}}$
whenever $a_{ij}+b_{ij}>0$ and $a_{ji}+b_{ji}>0$. Therefore, we have
$\sum_{i,j\in[n]\hskip 2.84526pt\textrm{s.t.}\hskip 2.84526pta_{ij}+b_{ij}>0}\frac{a_{ij}b_{ij}}{a_{ij}+b_{ij}}d_{X}(x_{i},x_{j})^{2}\leq\frac{1}{2}\sum_{i,j\in[n]\hskip 2.84526pt\textrm{s.t.}\hskip 2.84526pt\pi_{ij}+\pi_{ji}>0}\frac{\pi_{ij}\pi_{ji}}{\pi_{ij}+\pi_{ji}}d_{X}(x_{i},x_{j})^{2}.$
Combining this with the hypothesis that the inequality (6) holds true, we obtain the desired inequality. ∎
The following corollary is an immediate consequence of Lemma 2.11 and Corollary 2.10.
Corollary 2.12.
Let $n_{0}$ be a positive integer, and let $n\in[n_{0}]$. Let $X$ be a semimetric space that satisfies the $\mathrm{ANN}(n_{0})$ inequalities. Then any $x_{1},\ldots,x_{n}\in X$ satisfy the inequality (2) for any $m$, $c_{1},\ldots,c_{m}$, $\left(p_{i}^{1}\right)_{i\in[n]},\ldots,\left(p_{i}^{m}\right)_{i\in[n]}$, $\left(q_{j}^{1}\right)_{j\in[n]},\ldots,\left(q_{j}^{m}\right)_{j\in[n]}$, $A_{1},\ldots,A_{m}$, $B_{1},\ldots,B_{m}$ as in the statement of Theorem 1.4.
3. Lebedeva’s 6-point metric space
In this section, we recall the 6-point metric space constructed by Nina Lebedeva, which was appeared in [1, §7.2].
For any two points $a$ and $b$ in the Euclidean space $\mathbb{R}^{3}$, we denote by $[a,b]$ the line segment joining $a$ and $b$, and by $(a,b)$ the set $[a,b]\setminus\{a,b\}$. We also denote by $\mathrm{conv}(S)$ the convexhull of a subset $S$ of $\mathbb{R}^{3}$. Suppose $x_{1},\ldots,x_{6}$ are distinct six points in $\mathbb{R}^{3}$ that satisfy the following conditions:
$|(x_{1},x_{3})\cap(x_{2},x_{4})|$ $=1;$ (7)
$\left|(x_{5},x_{6})\cap\left(\mathrm{conv}(\{x_{1},x_{2},x_{3},x_{4}\})\setminus\cup_{\{i,j\}\in\binom{[4]}{2}}[x_{i},x_{j}]\right)\right|$ $=1.$ (8)
It follows from the conditions (7) and (8) that the points $x_{1},\ldots,x_{4}$ form the vertices of the convex quadrilateral in a $2$-dimensional affine subspace of $\mathbb{R}^{3}$, and that the points $x_{1},\ldots,x_{6}$ form the vertices of a nonregular convex octahedron in $\mathbb{R}^{3}$. We set $L=\{x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}\}$. For any $\varepsilon\geq 0$, we define a map $d_{\varepsilon}:L\times L\to[0,\infty)$ by
$d_{\varepsilon}(a,b)=\begin{cases}\|a-b\|+\varepsilon,\quad\textrm{if }\{a,b\}=\{x_{5},x_{6}\},\\ \|a-b\|,\quad\textrm{if }\{a,b\}\neq\{x_{5},x_{6}\},\end{cases}$ (9)
where $\|a-b\|$ is the Euclidean norm of $a-b$.
Theorem 3.1 (Lebedeva (see §7.2 of [1])).
Fix distinct six points $x_{1},\ldots,x_{6}\in\mathbb{R}^{3}$ that satisfy (7) and (8). Set $L=\{x_{1},\ldots,x_{6}\}$. Then there exists $C\in(0,\infty)$ such that for any $\varepsilon\in(0,C]$, $(L,d_{\varepsilon})$ defined by (9) is a metric space that satisfies the $(2+2)$-point comparison and the $(4+2)$-point comparison, but does not admit an isometric embedding into any $\mathrm{CAT}(0)$ space.
For the definition of the $(2+2)$-point comparison and the $(4+2)$-point comparison, see [1, §6.2]. It is known that a metric space satisfies the $(2+2)$-comparison if and only if it satisfies the $\boxtimes$-inequalities.
4. Proof of Theorem 1.5
In this section, we prove the following theorem, which implies Theorem 1.5. For any subset $M$ of $\mathbb{R}^{3}$ and any $R\in(0,\infty)$, we denote by $B(M,R)$ the open $R$-neighborhood of $M$ in $\mathbb{R}^{3}$. For any $x\in X$, we denote $B(\{x\},R)$ simply by $B(x,R)$.
Theorem 4.1.
Fix distinct six points $x_{1},\ldots,x_{6}\in\mathbb{R}^{3}$ that satisfy (7) and (8). Set $L=\{x_{1},\ldots,x_{6}\}$. Then there exists $C\in(0,\infty)$ such that for any $\varepsilon\in(0,C]$, $(L,d_{\varepsilon})$ defined by (9) satisfies the $\mathrm{ANN}(n)$ inequalities for all positive integers $n$.
Proof.
We first define four positive real constants $h$, $H$, $\theta$ and $\delta$ that depend only on the choice of $x_{1},\ldots,x_{6}\in\mathbb{R}^{3}$. We set
$h$ $=\min\left\{\min_{y\in\mathrm{conv}\left(L\setminus\left\{x_{5}\right\}\right)}\|x_{5}-y\|,\quad\min_{y\in\mathrm{conv}\left(L\setminus\left\{x_{6}\right\}\right)}\|x_{6}-y\|\right\},$
$H$ $=\max_{y\in\mathrm{conv}\left(L\setminus\left\{x_{6}\right\}\right)}\|x_{6}-y\|,\quad\theta=\min_{k\in[4]}\measuredangle x_{k}x_{6}x_{5},$
where we denote by $\measuredangle x_{k}x_{6}x_{5}\in[0,\pi]$ the interior angle measure at $x_{6}$ of the triangle with vertices $x_{k}$, $x_{6}$ and $x_{5}$. It follows from (7) and (8) that $h>0$, $H>0$ and $\theta>0$. It also follows from (7) and (8) that we can choose $\delta\in(0,\infty)$ such that
$B([x_{5},x_{6}],\delta)\cap B([x_{i},x_{j}],\delta)$ ==$\emptyset$,= ∅ ,
$B([x_{5},x_{6}],\delta)\cap B([x_{5},x_{i}],\delta)$ $\subseteq B\left(x_{5},\frac{h}{2}\right),\quad B([x_{5},x_{6}],\delta)\cap B([x_{6},x_{i}],\delta)\subseteq B\left(x_{6},\frac{h}{2}\right)$
for any $i,j\in[4]$.
Fix nonnegative real numbers $p_{1},\ldots,p_{6},q_{1},\ldots,q_{6}$ and a $6\times 6$ matrix $(\pi_{ij})$ with nonnegative real entries that satisfy
$\sum_{k=1}^{6}p_{k}=\sum_{k=1}^{6}q_{k}=1,\quad\sum_{k=1}^{6}\pi_{ik}=p_{i}+q_{i}$ (10)
for every $i\in[6]$. If $\pi_{56}=0$ or $\pi_{65}=0$, then the desired inequality
$\frac{1}{2}\sum_{i,j\in[6]\hskip 2.84526pt\textrm{s.t.}\hskip 2.84526pt\pi_{ij}+\pi_{ji}>0}\frac{\pi_{ij}\pi_{ji}}{\pi_{ij}+\pi_{ji}}d_{\varepsilon}(x_{i},x_{j})^{2}\leq\sum_{i=1}^{6}\sum_{j=1}^{6}p_{i}q_{j}d_{\varepsilon}(x_{i},x_{j})^{2}$
holds true for every $\varepsilon>0$ by definition of $d_{\varepsilon}$. So henceforth we assume that $\pi_{56}>0$ and $\pi_{65}>0$. We set
$t_{i}=\frac{\pi_{i6}}{\pi_{i6}+\pi_{6i}}.$
for each $i\in[5]$ with $\pi_{i6}+\pi_{6i}>0$.
For each $\varepsilon\geq 0$, we define $F(\varepsilon)\in\mathbb{R}$ by
[…]
Then it follows from the definition of $d_{\varepsilon}$ that
$F(\varepsilon)\geq\left(p_{5}q_{6}+p_{6}q_{5}-\frac{\pi_{56}\pi_{65}}{\pi_{56}+\pi_{65}}\right)f(\varepsilon)+F(0),$ (11)
where $f:[0,\infty)\to[0,\infty)$ is the strictly increasing function defined by
$f(\varepsilon)=2\|x_{5}-x_{6}\|\varepsilon+\varepsilon^{2},\quad\varepsilon\in[0,\infty).$
We define $z_{p}\in\mathbb{R}^{3}$, $z_{q}\in\mathbb{R}^{3}$ and $z_{ij}\in\mathbb{R}^{3}$ for any $i,j\in[6]$ with $\pi_{ij}+\pi_{ji}>0$ by
$z_{p}=\sum_{k=1}^{6}p_{k}x_{k},\quad z_{q}=\sum_{k=1}^{6}q_{k}x_{k},\quad z_{ij}=\frac{\pi_{ij}x_{i}+\pi_{ji}x_{j}}{\pi_{ij}+\pi_{ji}}.$
In other words, $z_{p}$, $z_{q}$ and $z_{ij}$ are the barycenters of the probability measures $\sum_{k=1}^{6}p_{k}\delta_{x_{k}}$, $\sum_{k=1}^{6}q_{k}\delta_{x_{k}}$ and $(\pi_{ij}\delta_{x_{i}}+\pi_{ji}\delta_{x_{j}})/(\pi_{ij}+\pi_{ji})$, respectively. It follows from Proposition 2.3 that
[…]
Combining these inequalities with (11), we obtain
$F(\varepsilon)\geq\left(p_{5}q_{6}+p_{6}q_{5}-\frac{\pi_{56}\pi_{65}}{\pi_{56}+\pi_{65}}\right)f(\varepsilon)\\ +\max\left\{\sum_{i,j\in[6]\hskip 2.84526pt\textrm{s.t. }\pi_{ij}+\pi_{ji}>0}\pi_{ij}\|z_{p}-z_{ij}\|^{2},\sum_{i,j\in[6]\hskip 2.84526pt\textrm{s.t. }\pi_{ij}+\pi_{ji}>0}\pi_{ij}\|z_{q}-z_{ij}\|^{2}\right\}.$ (12)
Set $S=\sum_{i=1}^{4}\sum_{j=1}^{6}\pi_{ij}$. Then we have
$p_{6}\geq 1-p_{5}-S,\quad q_{6}\geq 1-q_{5}-S,\quad\pi_{56}\leq p_{5}+q_{5},\quad\pi_{65}\leq 2-p_{5}-q_{5}.$
It follows that
$p_{5}q_{6}+p_{6}q_{5}-\frac{\pi_{56}\pi_{65}}{\pi_{56}+\pi_{65}}$ $\geq p_{5}(1-q_{5}-S)+(1-p_{5}-S)q_{5}-\frac{(p_{5}+q_{5})(2-p_{5}-q_{5})}{2}$
$=\frac{(p_{5}-q_{5})^{2}}{2}-(p_{5}+q_{5})S\geq-(p_{5}+q_{5})S$ (13)
since the function $f(x,y)=xy/(x+y)$, $x,y\in(0,\infty)$ is increasing with respect to both $x$ and $y$.
We will prove that there exists a constant $C>0$ such that $C$ depends only on the choice of the points $x_{1},\ldots,x_{6}$ in $\mathbb{R}^{3}$ and the inequality $F(\varepsilon)\geq 0$ holds for all $\varepsilon\in[0,C]$. We consider several cases separately. We first consider the following two cases.
Case 1: The inequality $\|z_{p}-z_{56}\|\geq\delta$ holds, or the inequality $\|z_{q}-z_{56}\|\geq\delta$ holds. In this case, it follows from (12) that
$F(\varepsilon)\geq-\frac{\pi_{56}\pi_{65}}{\pi_{56}+\pi_{65}}f(\varepsilon)+(\pi_{56}+\pi_{65})\delta^{2}\geq(\pi_{56}+\pi_{65})(\delta^{2}-f(\varepsilon)).$
Therefore, we have $F(\varepsilon)\geq 0$ for any $\varepsilon\in(0,f^{-1}(\delta^{2})]$.
Case 2: The inequality $\|z_{p}-z_{ij}\|\geq\delta$ holds for all $i,j\in[6]$ with $\{i,j\}\not\subseteq\{5,6\}$ and $\pi_{ij}+\pi_{ji}>0$, or the inequality $\|z_{q}-z_{ij}\|\geq\delta$ holds for all $i,j\in[6]$ with $\{i,j\}\not\subseteq\{5,6\}$ and $\pi_{ij}+\pi_{ji}>0$. In this case, it follows from (12) and (13) that
$F(\varepsilon)\geq-(p_{5}+q_{5})Sf(\varepsilon)+S\delta^{2}\geq S\left(\delta^{2}-2f(\varepsilon)\right).$
Therefore, we have $F(\varepsilon)\geq 0$ for any $\varepsilon\in(0,f^{-1}(\delta^{2}/2)]$.
Henceforth we assume that it is neither Case 1 nor Case 2. Then it follows from the definition of $\delta$ that we have
$\{z_{p},z_{q}\}\subseteq B\left(x_{5},\frac{h}{2}\right)\cup B\left(x_{6},\frac{h}{2}\right),$
which implies that
$\|z_{p}-z_{ij}\|>\frac{h}{2},\quad\|z_{q}-z_{ij}\|>\frac{h}{2}$ (14)
for any $i,j\in[4]$ with $\pi_{ij}+\pi_{ji}>0$.
We define $K_{0}\in[0,\infty)$ to be the nonnegative real number that satisfies
$\min\left\{p_{6},q_{6}\right\}=1-K_{0}t_{5}.$
Then we have
$p_{5}q_{6}+p_{6}q_{5}-\frac{\pi_{56}\pi_{65}}{\pi_{56}+\pi_{65}}$ $\geq(p_{5}+q_{5})\min\left\{p_{6},q_{6}\right\}-t_{5}(1-t_{5})(\pi_{56}+\pi_{65})$
$\geq\pi_{56}(1-K_{0}t_{5})-t_{5}(1-t_{5})(\pi_{56}+\pi_{65})=(\pi_{56}+\pi_{65})(1-K_{0})t_{5}^{2},$ (15)
which implies that the inequality $F(\varepsilon)\geq 0$ holds for any $\varepsilon>0$ whenever $K_{0}\leq 1$. So henceforth we assume that $K_{0}>1$. We define $z\in\mathbb{R}^{3}$ and $\tilde{z}\in\mathbb{R}^{3}$ by
$z=\begin{cases}z_{p},\quad\textrm{if }p_{6}\leq q_{6},\\ z_{q},\quad\textrm{if }q_{6}<p_{6},\end{cases}\quad\tilde{z}=\begin{cases}\frac{1}{1-p_{6}}\sum_{k=1}^{5}p_{k}x_{k},\quad\textrm{if }p_{6}\leq q_{6},\\ \frac{1}{1-q_{6}}\sum_{k=1}^{5}q_{k}x_{k},\quad\textrm{if }q_{6}<p_{6}.\end{cases}$
To complete the proof, we fix a constant $\upgamma\in(0,\infty)$, and consider three cases.
Case 3: The inequality $K_{0}\geq(1+\upgamma)\|x_{5}-x_{6}\|/h$ holds. In this case, we have
$\|z-z_{56}\|\geq\|z-x_{6}\|-\|z_{56}-x_{6}\|=K_{0}t_{5}\left\|\tilde{z}-x_{6}\right\|-t_{5}\|x_{5}-x_{6}\|\geq\left(K_{0}h-\|x_{5}-x_{6}\|\right)t_{5}>0.$
Together with (12) and (15), this implies that
$F(\varepsilon)>-(\pi_{56}+\pi_{65})(K_{0}-1)t_{5}^{2}f(\varepsilon)+(\pi_{56}+\pi_{65})\left(K_{0}h-\|x_{5}-x_{6}\|\right)^{2}t_{5}^{2}.$
It follows that we have $F(\varepsilon)>0$ whenever
$f(\varepsilon)\leq\frac{(K_{0}h-\|x_{5}-x_{6}\|)^{2}}{K_{0}-1}.$
Since $\|x_{5}-x_{6}\|/h\geq 1$, it is easily seen that the function
$\varphi_{0}(K)=\frac{(Kh-\|x_{5}-x_{6}\|)^{2}}{K-1},\quad K\in\left[\frac{(1+\upgamma)\|x_{5}-x_{6}\|}{h},\infty\right)$
is increasing. Therefore, we have $F(\varepsilon)>0$ whenever
$\varepsilon\leq f^{-1}\left(\frac{h\upgamma^{2}\|x_{5}-x_{6}\|^{2}}{(1+\upgamma)\|x_{5}-x_{6}\|-h}\right).$
Case 4: The inequalities $1<K_{0}<(1+\upgamma)\|x_{5}-x_{6}\|/h$ and $\measuredangle zx_{6}x_{5}\geq\theta/2$ hold. In this case, we have
$\|z-z_{56}\|\geq\left(\sin\frac{\theta}{2}\right)\|x_{5}-x_{6}\|t_{5}.$
Together with (12) and (15), this implies that
$F(\varepsilon)>(\pi_{56}+\pi_{65})\left(-\left(\frac{(1+\upgamma)\|x_{5}-x_{6}\|}{h}-1\right)f(\varepsilon)+\left(\sin\frac{\theta}{2}\right)^{2}\|x_{5}-x_{6}\|^{2}\right)t_{5}^{2}.$
Therefore, we have $F(\varepsilon)>0$ whenever
$\varepsilon\leq f^{-1}\left(\frac{h\left(\sin\frac{\theta}{2}\right)^{2}\|x_{5}-x_{6}\|^{2}}{(1+\upgamma)\|x_{5}-x_{6}\|-h}\right).$
Case 5: The inequalities $1<K_{0}<(1+\upgamma)\|x_{5}-x_{6}\|/h$ and $\measuredangle zx_{6}x_{5}<\theta/2$ hold. In this case, we have $\measuredangle zx_{6}x_{k}>\theta/2$ for any $k\in[4]$. Therefore we have
$\|z-z_{k6}\|>\left(\sin\frac{\theta}{2}\right)\|z-x_{6}\|=\left(\sin\frac{\theta}{2}\right)\|\tilde{z}-x_{6}\|K_{0}t_{5}>\left(\sin\frac{\theta}{2}\right)ht_{5}$ (16)
for any $k\in[4]$. Suppose $l$ and $m$ are positive integers at most $4$ such that
$\pi_{l6}>0,\quad\pi_{m6}>0,\quad t_{l}\|x_{l}-x_{6}\|\leq\frac{(1+\upgamma)^{2}\|x_{5}-x_{6}\|Ht_{5}}{h},\quad t_{m}\|x_{m}-x_{6}\|>\frac{(1+\upgamma)^{2}\|x_{5}-x_{6}\|Ht_{5}}{h}.$
Then we have
$0<t_{l}\leq\frac{(1+\upgamma)^{2}\|x_{5}-x_{6}\|Ht_{5}}{h^{2}}.$
Together with (16), this implies that
$(\pi_{l6}+\pi_{6l})\|z-z_{l6}\|^{2}=\frac{1}{t_{l}}\pi_{l6}\|z-z_{l6}\|^{2}>\frac{h^{4}\left(\sin\frac{\theta}{2}\right)^{2}}{(1+\upgamma)^{2}\|x_{5}-x_{6}\|H}\pi_{l6}t_{5}.$ (17)
We also have
$\|z-z_{m6}\|\geq\|z_{m6}-x_{6}\|-\|z-x_{6}\|=t_{m}\|x_{m}-x_{6}\|-K_{0}t_{5}\|\tilde{z}-x_{6}\|\\ >t_{m}\|x_{m}-x_{6}\|-\frac{(1+\upgamma)\|x_{5}-x_{6}\|Ht_{5}}{h}>\frac{\upgamma(1+\upgamma)\|x_{5}-x_{6}\|Ht_{5}}{h}>0.$
We define $K_{1}\in(\upgamma(1+\upgamma)\|x_{5}-x_{6}\|H/h,\infty)$ to be the positive real number that satisfies
$K_{1}t_{5}=t_{m}\|x_{m}-x_{6}\|-\frac{(1+\upgamma)\|x_{5}-x_{6}\|Ht_{5}}{h}.$
Then we have
$\|z-z_{m6}\|>K_{1}t_{5},\quad\pi_{m6}+\pi_{6m}=\frac{1}{t_{m}}\pi_{m6}\geq\frac{h^{2}}{\left(hK_{1}+(1+\upgamma)\|x_{5}-x_{6}\|H\right)t_{5}}\pi_{m6},$
and therefore
$(\pi_{m6}+\pi_{6m})\|z-z_{m6}\|^{2}>\frac{h^{2}K_{1}^{2}}{hK_{1}+(1+\upgamma)\|x_{5}-x_{6}\|H}\pi_{m6}t_{5}.$
Because the function
$\varphi_{1}(K)=\frac{h^{2}K^{2}}{hK+(1+\upgamma)\|x_{5}-x_{6}\|H},\quad K\in\left[\frac{\upgamma(1+\upgamma)\|x_{5}-x_{6}\|H}{h},\infty\right)$
is increasing, we obtain
$(\pi_{m6}+\pi_{6m})\|z-z_{m6}\|^{2}>\upgamma^{2}\|x_{5}-x_{6}\|H\pi_{m6}t_{5}.$ (18)
It follows from (14) that we have
$\|z-z_{ij}\|^{2}>\frac{h^{2}}{4}\geq\frac{h^{2}}{4}t_{5}$ (19)
for any $i,j\in[4]$. Set
$c=\min\left\{\frac{h^{4}\left(\sin\frac{\theta}{2}\right)^{2}}{(1+\upgamma)^{2}\|x_{5}-x_{6}\|H},\quad\upgamma^{2}\|x_{5}-x_{6}\|H,\quad\frac{h^{2}}{4}\right\}.$
Then it follows from (17), (18) and (19) that
$\sum_{i,j\in[n]\hskip 2.84526pt\textrm{s.t.}\pi_{ij}+\pi_{ji}>0}\pi_{ij}\|z-z_{ij}\|^{2}>cSt_{5}.$ (20)
On the other hand, it follows from (13) that
$p_{5}q_{6}+q_{5}p_{6}-\frac{\pi_{56}\pi_{65}}{\pi_{56}+\pi_{65}}\geq-(p_{5}+q_{5})S\\ \geq-(2-p_{6}-q_{6})S\geq-2K_{0}St_{5}>-\frac{2(1+\upgamma)\|x_{5}-x_{6}\|}{h}St_{5}.$ (21)
It follows from (12), (20) and (21) that
$F(\varepsilon)\geq-\frac{2(1+\upgamma)\|x_{5}-x_{6}\|}{h}St_{5}f(\varepsilon)+cSt_{5}.$
Therefore we have $F(\varepsilon)\geq 0$ whenever
$\varepsilon\leq f^{-1}\left(\frac{ch}{2(1+\upgamma)\|x_{5}-x_{6}\|}\right).$
So far, we have proved that there exists a constant $C>0$ such that $C$ depends only on the choice of the points $x_{1},\ldots,x_{6}\in\mathbb{R}^{3}$, and any $\varepsilon\in[0,C]$ satisfies
$\frac{1}{2}\sum_{i,j\in[6]\hskip 2.84526pt\textrm{s.t.}\hskip 2.84526pt\pi_{ij}+\pi_{ji}>0}\frac{\pi_{ij}\pi_{ji}}{\pi_{ij}+\pi_{ji}}d_{\varepsilon}(x_{i},x_{j})^{2}\leq\sum_{i=1}^{6}\sum_{j=1}^{6}p_{i}q_{j}d_{\varepsilon}(x_{i},x_{j})^{2}$
for any nonnegative real numbers $p_{1},\ldots,p_{6},q_{1},\ldots,q_{6}$ and any $6\times 6$ matrix $(\pi_{ij})$ with nonnegative entries that satisfy (10). Therefore, it follows from Proposition 2.9 that $(L,d_{\varepsilon})$ satisfies the $\mathrm{ANN}(n)$ inequalities for all positive integers $n$ whenever $\varepsilon\in[0,C]$. ∎
Theorem 1.5 and Corollary 1.6 follow from Theorem 4.1 immediately.
Proof of Theorem 1.5.
Theorem 3.1 and Theorem 4.1 imply Theorem 1.5. ∎
Proof of Corollary 1.6.
Theorem 1.5 and Corollary 2.12 imply Corollary 1.6. ∎

### v2

@@
 For any two points $a$ and $b$ in the Euclidean space $\mathbb{R}^{3}$, we denote by $[a,b]$ the line segment joining $a$ and $b$, and by $(a,b)$ the set $[a,b]\setminus\{a,b\}$.
-We also denote by $\mathrm{conv}(S)$ the convexhull of a subset $S$ of $\mathbb{R}^{3}$.
+We denote by $\mathrm{conv}(S)$ the convexhull of a subset $S$ of $\mathbb{R}^{3}$.
 Suppose $x_{1},\ldots,x_{6}$ are distinct six points in $\mathbb{R}^{3}$ that satisfy the following conditions:
@@
 $h$ $=\min\left\{\min_{y\in\mathrm{conv}\left(L\setminus\left\{x_{5}\right\}\right)}\|x_{5}-y\|,\quad\min_{y\in\mathrm{conv}\left(L\setminus\left\{x_{6}\right\}\right)}\|x_{6}-y\|\right\},$
-$H$ $=\max_{y\in\mathrm{conv}\left(L\setminus\left\{x_{6}\right\}\right)}\|x_{6}-y\|,\quad\theta=\min_{k\in[4]}\measuredangle x_{k}x_{6}x_{5},$
-where we denote by $\measuredangle x_{k}x_{6}x_{5}\in[0,\pi]$ the interior angle measure at $x_{6}$ of the triangle with vertices $x_{k}$, $x_{6}$ and $x_{5}$.
+$H$ $=\max\left\{\max_{y\in\mathrm{conv}\left(L\setminus\left\{x_{5}\right\}\right)}\|x_{5}-y\|,\quad\max_{y\in\mathrm{conv}\left(L\setminus\left\{x_{6}\right\}\right)}\|x_{6}-y\|\right\},$
+$\theta$ $=\min\left\{\min_{k\in[4]}\measuredangle x_{k}x_{5}x_{6},\quad\min_{k\in[4]}\measuredangle x_{k}x_{6}x_{5}\right\}$
+where we denote by $\measuredangle x_{k}x_{i}x_{j}\in[0,\pi]$ the interior angle measure at $x_{i}$ of the triangle with vertices $x_{k}$, $x_{i}$ and $x_{j}$.
 It follows from (7) and (8) that $h>0$, $H>0$ and $\theta>0$.
-It also follows from (7) and (8) that we can choose $\delta\in(0,\infty)$ such that
+It also follows from (7) and (8) that we can choose $\delta\in(0,h/2)$ such that we have
 $B([x_{5},x_{6}],\delta)\cap B([x_{i},x_{j}],\delta)$ ==$\emptyset$,= ∅ ,
@@
 $\frac{1}{2}\sum_{i,j\in[6]\hskip 2.84526pt\textrm{s.t.}\hskip 2.84526pt\pi_{ij}+\pi_{ji}>0}\frac{\pi_{ij}\pi_{ji}}{\pi_{ij}+\pi_{ji}}d_{\varepsilon}(x_{i},x_{j})^{2}\leq\sum_{i=1}^{6}\sum_{j=1}^{6}p_{i}q_{j}d_{\varepsilon}(x_{i},x_{j})^{2}$
-holds true for every $\varepsilon>0$ by definition of $d_{\varepsilon}$. So henceforth we assume that $\pi_{56}>0$ and $\pi_{65}>0$. We set
-$t_{i}=\frac{\pi_{i6}}{\pi_{i6}+\pi_{6i}}.$
-for each $i\in[5]$ with $\pi_{i6}+\pi_{6i}>0$.
+holds true for every $\varepsilon>0$ by definition of $d_{\varepsilon}$. So henceforth we assume that $\pi_{56}>0$ and $\pi_{65}>0$.
 For each $\varepsilon\geq 0$, we define $F(\varepsilon)\in\mathbb{R}$ by
@@
 Henceforth we assume that it is neither Case 1 nor Case 2. Then it follows from the definition of $\delta$ that we have
-$\{z_{p},z_{q}\}\subseteq B\left(x_{5},\frac{h}{2}\right)\cup B\left(x_{6},\frac{h}{2}\right),$
-which implies that
-$\|z_{p}-z_{ij}\|>\frac{h}{2},\quad\|z_{q}-z_{ij}\|>\frac{h}{2}$ (14)
-for any $i,j\in[4]$ with $\pi_{ij}+\pi_{ji}>0$.
-We define $K_{0}\in[0,\infty)$ to be the nonnegative real number that satisfies
+$\{z_{p},z_{q}\}\subseteq B\left(x_{5},\frac{h}{2}\right)$ (14)
+or
+$\{z_{p},z_{q}\}\subseteq B\left(x_{6},\frac{h}{2}\right).$ (15)
+Because it is easily seen that we can obtain the same estimate of $F(\varepsilon)$ in the same way as we describe below whether (14) holds or (15) holds, we assume that (15) holds.
+We set
+$t_{i}=\frac{\pi_{i6}}{\pi_{i6}+\pi_{6i}}.$
+for each $i\in[5]$ with $\pi_{i6}+\pi_{6i}>0$, and define $K_{0}\in[0,\infty)$ to be the nonnegative real number that satisfies
 $\min\left\{p_{6},q_{6}\right\}=1-K_{0}t_{5}.$
@@
 It follows from (15) that we have
-$\|z-z_{ij}\|^{2}>\frac{h^{2}}{4}\geq\frac{h^{2}}{4}t_{5}$ (19)
-for any $i,j\in[4]$. Set
+$\|z_{p}-z_{ij}\|>\frac{h}{2},\quad\|z_{q}-z_{ij}\|>\frac{h}{2}$
+for any $i\in[4]$ and any $j\in[5]$. Therefore, we have
+$\pi_{ij}\|z-z_{ij}\|^{2}>\frac{h^{2}}{4}\pi_{ij}\geq\frac{h^{2}}{4}\pi_{ij}t_{5}$ (20)
+for any $i\in[4]$ and any $j\in[5]$. Set
 $c=\min\left\{\frac{h^{4}\left(\sin\frac{\theta}{2}\right)^{2}}{(1+\upgamma)^{2}\|x_{5}-x_{6}\|H},\quad\upgamma^{2}\|x_{5}-x_{6}\|H,\quad\frac{h^{2}}{4}\right\}.$